# Bending-strain effects in conventional superconductors and superconducting junctions

### v1

@@ I Introduction
Modulation of the singlet order parameter occurs under the effect of SOC, an external magnetic field or an internal exchange field. When a superconductor is proximity-coupled to a ferromagnet or antiferromagnet, or exposed to an external magnetic field, odd-frequency, s-wave spin-polarized triplet correlations can be generated. That is, the generation of these correlations requires both spin-splitting and a second spin-orienting field, typically engineered via magnetic multilayers Eschrig (2011); Robinson et al. (2010); Khaire et al. (2010) or by employing intrinsic SOC (for example from a heavy metal) in combination with a single exchange field Bergeret and Tokatly (2013, 2014); Jacobsen et al. (2015). Recent studies have shown that geometric curvature, increasingly harnessed for its advantages in spintronics Gentile et al. (2022); Ortix (2015); Gentile et al. (2013); Ying et al. (2016); Das et al. (2019); Makarov et al. (2022); Fomin and Dobrovolskiy (2022); Kopasov et al. (2021), is an effective way to both design and manipulate curvature-induced SOC. Geometric curvature alone is therefore sufficient to affect the symmetries of superconducting correlations persisting in single magnetic elements i.e. without using any intrinsic SOC, magnetic multilayers or strain Salamone et al. (2021, 2022); Skarpeid et al. (2024); Salamone et al. (2024a, b). The addition of bending-strain may then be used to manipulate the different superconductivity populations in-situ, for example to control the superconducting transition Salamone et al. (2022, 2024b), or the direction of Josephson current Salamone et al. (2021); Skarpeid et al. (2024).
@@
In contrast to its role in magnetic materials, geometric curvature alone (i.e. curvature without strain) is not expected to affect the symmetries of the order parameter of a plain s-wave superconductor. With ballistic transport, an interface with a normal metal (N) will convert a portion of the even-frequency, s-wave singlet correlations into odd-frequency p-wave singlets Tanaka et al. (2007). When proximity-coupled to a non-magnetic material such as a semiconductor, normal metal or topological insulator, geometric curvature in the non-magnetic material can then affect the singlet order parameter in the presence of SOC, due to its tunable effect on the generation of unpolarised triplet correlations Ying et al. (2017); Kutlin and Mel’nikov (2020). For example, curvature in a proximity-coupled topological insulator, which has strong SOC, can induce topological superconductivity Chou et al. (2021). Spin-orbit-coupled superconducting wires have also been found to display the evolution of topological phases themselves due to geometric deformation Francica et al. (2020). Bending-strain introduces an effective SOC with a rotating normal vector. While of Rashba-type, the rotating angle affects the generation and modulation of triplet correlations differently from a constant, intrinsic Rashba-SOC, or extrinsic SOC from gating. In this article, we explore the extent to which bending-strain affects the symmetries of the order parameter in a conventional, even-frequency s-wave spin-singlet superconductor alone, and we examine the implications of the symmetry transformations on the proximity effect with normal metals in bilayers and Josephson junctions.
@@
We will use the Bogoliubov-de Gennes framework Zhu (2016) to examine how the effective SOC of bending-strain in clean, conventional superconductors (S) can modulate the symmetries of the order parameter, and we therefore begin by introducing the curvilinear framework and discretizing and diagonalizing the continuum Hamiltonian in Section II. We go on to show how this combines with symmetry-modulations from normal-metal interfaces in SN bilayers and SNS Josephson junctions in Section III. We find that bending-strain in the superconductor gives an effective SOC that generates even-frequency p-wave triplets, both with and without spin polarization. In bilayers, the magnitude of the additional odd-frequency pairings can be tuned with the strain, and we find an in-plane, strain-induced magnetization that switches sign as a function of the strain in SNS junctions. We conclude by discussing the potential benefits for superconducting spintronic devices in Section IV.
II Theory
We introduce geometric curvature in superconductors using the Frenet-Serret frame in Section II.1, and define the continuum Hamiltonian with curvature-induced bending-strain in Section II.2. We discretize the Hamiltonian to give the tight-binding model in Section II.3. In Section II.4, we define the s- and p-wave singlet and triplet amplitudes, as well as the density of states and spin magnetization as observables.
II.1 Frenet-Serret frame
@@
Here we have introduced the geometric curvature function $\kappa(s)$, which for a uniform curve is simply the inverse of the radius of the curve (see Fig. 1). Using Eq. 1, we can write the basis vectors as derivatives of $\bm{R}(s,n,b)$. We express them as $\bm{e}_{\mu}=\partial_{\mu}\bm{R}=h_{\mu}\hat{e}_{\mu}$, where $h_{\mu}$ are scale factors.
Figure 1: Thin-film superconductor with bending-strain and an SNS Josephson junction with uniformly curved superconductors and an uncurved normal metal. Curvilinear unit vectors in the tangential $\hat{\mathcal{T}}(s)$, normal $\hat{\mathcal{N}}(s)$, and binormal ^$\hat{\mathcal{B}}(s)$ direction, as well as the radius of curvature $R$, are indicated. In the inset, red represents tensile strain and blue indicates compressive strain of the lattice.
The metric tensor is defined as ${\mathcal{G}_{\lambda\mu}=\bm{e}_{\lambda}\cdot\bm{e}_{\mu}}$, and can be written as
@@
$\bm{r}(s)=\frac{1}{\kappa}\cos(\kappa s)\>\hat{\bm{e}}_{x}+\frac{1}{\kappa}\sin(\kappa s)\>\hat{\bm{e}}_{y}.$ (3)
The orthogonal curvilinear unit vectors become $\hat{\mathcal{T}}(s)=\partial_{s}\bm{r}(s)$, $\hat{\mathcal{N}}(s)=\partial_{s}\hat{\mathcal{T}}(s)/|\partial_{s}\hat{\mathcal{T}}(s)|$ and ^$\hat{\mathcal{B}}(s)=\hat{\mathcal{T}}(s)\times\hat{\mathcal{N}}(s)$. For the circular parametrization [Eq. 3] this gives
$\begin{split}\hat{\mathcal{T}}(s)&=-\sin(\kappa s)\>\hat{\bm{e}}_{x}+\cos(\kappa s)\>\hat{\bm{e}}_{y},\\ \hat{\mathcal{N}}(s)&=-\cos(\kappa s)\>\hat{\bm{e}}_{x}-\sin(\kappa s)\>\hat{\bm{e}}_{y},\\ \hat{\mathcal{B}}(s)&=\hat{\bm{e}}_{z}.\end{split}$ (4)
II.2 Continuum Hamiltonian and strain-induced spin-orbit coupling

### v2

@@ I Introduction
Modulation of the singlet order parameter occurs under the effect of SOC, an external magnetic field or an internal exchange field. When a superconductor is proximity-coupled to a ferromagnet or antiferromagnet, or exposed to an external magnetic field, odd-frequency, s-wave spin-polarized triplet correlations can be generated. That is, the generation of these correlations requires both spin-splitting and a second spin-orienting field, typically engineered via magnetic multilayers or conical exchange magnets Eschrig (2011); Robinson et al. (2010); Khaire et al. (2010), or by employing intrinsic SOC (for example from a heavy metal) in combination with a single exchange field Bergeret and Tokatly (2013, 2014); Jacobsen et al. (2015). Recent studies have shown that geometric curvature, increasingly harnessed for its advantages in spintronics Gentile et al. (2022); Ortix (2015); Gentile et al. (2013); Ying et al. (2016); Das et al. (2019); Makarov et al. (2022); Fomin and Dobrovolskiy (2022); Kopasov et al. (2021), is an effective way to both design and manipulate curvature-induced SOC. Geometric curvature alone is therefore sufficient to affect the symmetries of superconducting correlations persisting in single magnetic elements i.e. without using any intrinsic SOC, magnetic multilayers or strain Salamone et al. (2021, 2022); Skarpeid et al. (2024); Salamone et al. (2024a, b). The addition of bending-strain may then be used to manipulate the different superconductivity populations in-situ, for example to control the superconducting transition Salamone et al. (2022, 2024b), or the direction of Josephson current Salamone et al. (2021); Skarpeid et al. (2024).
@@
In contrast to its role in magnetic materials, geometric curvature alone (i.e. curvature without strain) is not expected to affect the symmetries of the order parameter of a plain s-wave superconductor (S). With ballistic transport, an interface with a normal metal (N) will convert a portion of the even-frequency, s-wave singlet correlations into odd-frequency p-wave singlets Tanaka et al. (2007). When proximity-coupled to a non-magnetic material such as a semiconductor, normal metal or topological insulator, geometric curvature in the non-magnetic material can then affect the singlet order parameter in the presence of SOC, due to its tunable effect on the generation of unpolarised triplet correlations Ying et al. (2017); Kutlin and Mel’nikov (2020). For example, curvature in a proximity-coupled topological insulator, which has strong SOC, can induce topological superconductivity Chou et al. (2021). Spin-orbit-coupled superconducting wires have also been found to display the evolution of topological phases themselves due to geometric deformation Francica et al. (2020). Bending-strain introduces an effective SOC with a rotating normal vector. While of Rashba-type, the rotating angle affects the generation and modulation of triplet correlations differently from a constant, intrinsic Rashba-SOC, or extrinsic SOC from gating. In this article, we explore the extent to which bending-strain affects the symmetries of the order parameter in a conventional, even-frequency s-wave spin-singlet superconductor alone, and we examine the implications of the symmetry transformations on the proximity effect with normal metals in bilayers and Josephson junctions.
@@
We will use the Bogoliubov-de Gennes framework Zhu (2016) to show how the effective SOC of bending-strain in clean, conventional superconductors (see Figure 1) can modulate the symmetries of the order parameter, generating even-frequency p-wave triplets, both with and without spin polarization. We will show how this combines with symmetry-modulations from normal-metal interfaces in SN bilayers and SNS Josephson junctions with a straight normal metal: In bilayers, the magnitude of the additional odd-frequency pairings can be tuned with the strain, and we find an in-plane, strain-induced magnetization that switches sign as a function of the strain in SNS junctions.
Figure 1: Thin-film superconductor with bending-strain along a constant curve with arclength $s$ (top), and (below) an SNS Josephson junction with uniformly curved superconductors and an uncurved normal metal. Orthonormal, curvilinear unit vectors in the tangential $\hat{\mathcal{T}}(s)$, normal $\hat{\mathcal{N}}(s)$, and binormal ^$\hat{\mathcal{B}}(s)$ direction, as well as the radius of curvature $R$, are indicated. In the inset, red represents tensile strain and blue indicates compressive strain of the lattice.
II Results
The order parameters and observables we will examine are defined and derived in curvilinear coordinates in Section II.1. In Section II.2, we explore the effects of bending-strain in clean, conventional superconducting films. In Section II.3, we consider the additional effect of proximity-coupling the strained superconductor to a normal metal, and in Section II.4 we consider the proximity-induced effects in SNS junctions. The theoretical details, including how strain-induced SOC enters in the curvilinear framework, as well as the discretization and diagonalization of the continuum Hamiltonian, are provided in the Methods, Section IV. Parameter regimes for the numerical studies are explained and illustrated in Section IV.4.
II.1 Order parameters and observables
To compare different superconducting symmetries, we define the order parameters as $\Delta_{\bm{i}}(\tau)=\langle\mathcal{T}c_{\bm{i},\sigma}(\tau)\>c_{\bm{i}^{\prime},\sigma^{\prime}}(0)\rangle$, where $\mathcal{T}$ is the time ordering operator, and $\tau$ is the relative time coordinate scaled by $\hbar/t$. The hopping amplitude between nearest neighbours is denoted by $t$, and $c_{\bm{i},\sigma}(\tau)$ are the fermion annihilation operators for spin $\sigma$ at lattice site $\bm{i}$ [see Methods Section IV.3, for details]. Thus, we can investigate the order parameters of even- ($\tau=0$) and odd-frequency ($\tau\neq 0$) correlations. In Table 1, we introduce the notation for the relevant combinations of underlying Spin-Parity-Orbital-Time symmetries that ensure overall odd order parameter symmetry Linder and Balatsky (2019).
-$S$$P^{*}$$O$$T^{*}$
$\mathcal{S}_{0}$1 -1 -1 -1 even-frequency $s$-wave singlet
$\mathcal{S}_{\sigma\sigma^{\prime}}(\tau)$1 -1 -1 -1 odd-frequency $s$-wave triplets
$\mathcal{P}^{\bm{m}}_{\sigma\sigma^{\prime}}$1 -1 -1 -1 even-frequency $p^{m}$-wave triplets
$\mathcal{P}^{\bm{m}}_{0}(\tau)$1 -1 -1 -1 odd-frequency $p^{m}$-wave singlet
Here $\bm{m}$ is one lattice spacing in either the tangential or binormal direction (see Fig. 1), and the time-dependent operators are given by $c_{\bm{i},\sigma}(\tau)=e^{i\mathcal{H}\tau}c_{\bm{i},\sigma}e^{-i\mathcal{H}\tau}$. Observe here that the p-wave correlations have $\bm{i}-\bm{i}^{\prime}\neq 0$, and the s-wave correlations have $\bm{i}-\bm{i}^{\prime}=0$.
The gap equation in the mean-field approximation is given by $\Delta_{i}=U_{i}\>\mathcal{S}_{i,0}$ and can be written as
where we have introduced a new quasiparticle basis $\gamma_{i,n,k}$ through a Bogoliubov transformation with complex coefficients $\{u_{i,n,k},v_{i,n,k},w_{i,n,k},x_{i,n,k}\}$ [details in Methods, Section IV]. Furthermore, we have used that ${\big{<}{\gamma^{\dagger}_{n,k}\gamma_{n^{\prime},k}}\big{>}=f(2\mathcal{E}_{n,k})\delta_{n,n^{\prime}}}$, where $\mathcal{E}_{n,k}$ is the eigenenergy associated with $\gamma_{i,n,k}$ and $f(2\mathcal{E}_{n,k})$ is the Fermi-Dirac distribution, and $\beta$ is the inverse temperature. The corresponding expressions for Eq. S1-(S5) in the new quasiparticle basis [Eq. 33] that are used for computation are given in Sec. S.I., in the Supplementary Information.
$\sum_{m}\bm{J}_{\bm{i},m}=-i\big{[}\mathcal{H},\>\bm{S}_{\bm{i}}\big{]}.$ (8)
Here, $\mathcal{H}$ is the Hamiltonian and $m$ enumerates the surface normals of the unit cell $\bm{i}$. By comparing the current flowing in and out from a single unit cell in the tangential direction, a final expression can be obtained Risinggård and Linder (2019). Thus, $\bm{J}$ denotes the spin-current flowing parallel to the tangential unit vector $\hat{\mathcal{T}}$ in real space, with components in spin space $J^{\mu}$, where $\mu=(s,n,b)$ indicates spin alignment in the tangential, normal, and binormal directions. The components of the spin-current $\bm{J}$ are only conserved in areas without spin-orbit coupling. We will therefore only examine spin-current as a useful observable in cases where the material is not curved at $\bm{i}$, i.e. in an uncurved normal metal sandwiched between two curved superconductors. Applying the standard quasiparticle operators and coefficients [Eq. 29 in Methods, Section IV], we find the tangential- and binormal-components of the spin-current
II.2 Bending-strain in a thin-film superconductor
The superconductor is initialized with an s-wave gap on a square lattice, and the gap equation is solved self-consistently by iterative diagonalization (details of the numerical implementation are provided in the Methods, Section IV.4). Bending the superconducting film creates a strain-induced spin-orbit coupling, where the direction of asymmetry is in the normal direction to the surface. This creates an additional term in the Hamiltonian described by
${\displaystyle\mathcal{H}_{\text{soc}}=\alpha_{N}\big{(}\sigma_{B}\hat{k}_{s}-\sigma_{T}\hat{k}_{b}\big{)},}$ (12)
where $\alpha_{N}$ is proportional to the local curvature, $\sigma_{\mu}$ are curvilinear Pauli-matrices, and $\hat{k}_{\mu}$ are momentum operators. This strain-induced SOC creates a mixed parity s- and p-wave order parameter in the superconductor. The magnitude of the s-wave gap will diminish as a function of the strain, as shown in Fig. 2. This is analogous to doing self-consistency calculations for an increasing Rashba coefficient in a spin-orbit coupled superconductor, which yields the same results Ying et al. (2017).
We denote the p-wave triplet pairings with no spin projection as $\mathcal{P}^{s}_{\uparrow\downarrow}$, and the spin-polarized triplets as $\mathcal{P}^{b}_{\sigma}$, as indicated in Table 1. With time-inversion symmetry conserved, the spin-polarized p-wave amplitudes are equal in magnitude $|\mathcal{P}^{b}_{\uparrow}|=|\mathcal{P}^{b}_{\downarrow}|$. The superconducting gap closes completely for high enough strain, so that both s- and p-wave correlations become negligible. The thickness of the film denoted as $N_{n}$, determines the maximum strain and the extent of superconductivity it can support. We assume that the strength of the SOC is proportional to the curvature, which means that a thinner film can accommodate greater curvature for the same surface strain. The curvature at a given strain is less for thicker films, and so the effective SOC is weaker. The singlet s-wave and finite triplet p-wave order pairings are shown in Fig. 2 as a function of the strain. While the curvature is in the tangential direction, the pairings $\mathcal{P}^{s}_{\uparrow\downarrow}$ and $\mathcal{P}^{b}_{\sigma}$ experience the same strain-induced electric field in the normal direction, leading to their similar profile in Fig. 2. Their spin polarization depends on the chosen spin axis [c.f. Eq. 22] and, in our example, both the spin axis and $\hat{\mathcal{B}}$ are aligned parallel to $\hat{z}$. The pairings arise through the entire superconductor because the superconductor has uniform circular curvature, giving a constant bending-strain profile and equivalent SOC-strength. If the strain is only applied to a certain part, the p-wave amplitudes will only be finite locally. For instance, in an ellipse, the correlations are localized at the apexesHeinrich (2024). The emergence of $\mathcal{P}^{s}$-wave correlations under bending strain aligns with existing intuition from one-dimensional systems. However, in thin films, the inclusion of the binormal direction and the term $\sigma_{T}\hat{k}_{b}$, which here leads to the emergence of $\mathcal{P}^{b}$-wave triplet states in strained superconducting films, is crucial to the appearance of novel effects in conjunction with interfacial symmetry breaking and Josephson currents, as we will now explore.
II.3 SN bilayer with bending-strain in S
We consider a superconductor-normal metal (SN) bilayer, with an example of arbitrarily chosen length of superconductor with $N_{S}=50$ sites, a normal metal with $N_{N}=21$ sites, and $N_{b}=71$ sites in the binormal direction. Choosing an odd number simplifies the numerical algorithm at the edge of the Brillouin zone, but does not impact the results (see Section IV.4). The film thickness in the normal direction is in this case set to $N_{n}=3$, and the strain is only applied to the superconductor.
In Fig. 3, we show the even-frequency pairing amplitudes and the relative time derivatives of the finite odd-frequency pairings at the interface. The even-frequency amplitudes originate from the bending-strain via the spin-asymmetry introduced by the effective spin-orbit coupling in the superconductor, while the interface between the materials induces the odd-frequency pairings. The even-frequency p-wave triplets $\mathcal{P}^{s}_{\uparrow\downarrow}$ persist throughout the bulk, and these triplets have an interface-dependence on both Andreev and normal reflections Cayao and Black-Schaffer (2018). They vanish for $\alpha_{N}=0$, in contrast to $\mathcal{P}^{s}_{0}(\tau)$, which remains finite, as can be seen in Fig. 3. Furthermore, since the superconductor is strained and the normal metal is not, the curvature-induced spin-orbit coupling vanishes at the interface. This interfacial change acts on spins like a local spin splitting in the binormal direction, giving rise to an odd-frequency $\mathcal{S}_{\uparrow\downarrow}(\tau)$-wave triplet and $\mathcal{P}^{b}_{0}(\tau)$-wave singlet contribution.
Figure 3: The pairing amplitudes around the SN interface for increasing strain in the superconductor. The first column shows the non-zero even-frequency pairings and the second column is the relative time derivative of the odd-frequency pairings at $\tau=0$. The spin-polarized p-wave pairings are equal in magnitude $|\mathcal{P}^{b}_{\sigma}|=|\mathcal{P}^{b}_{-\sigma}|$. The derivatives are compared to the amplitudes of the corresponding odd-frequency pairings in Fig. S.1 in the Supplementary Information. System parameters are given in the main text.
For completeness, we compare the derivatives in Fig. 3 with the magnitudes of the corresponding odd-frequency pairings in Fig. S.1 in the Supplementary Information. There we see that the amplitudes – which remain within the same order for relative times within that order – are comparable to the derivative evaluated at relative time $\tau=0$.
II.4 SNS junction with bending-strain in S
In Fig. 4, we show the tangential and binormal components of the spin current $\bm{J}$ at the center of the normal metal. The sign of $J^{b}$ depends on the length of the normal metal and strain, resulting in so-called $0-\pi$ oscillations as a function of both length and the effective spin-orbit coupling strength experienced, shown here as a function of the bending-strain in the superconductor. The $0-\pi$ oscillations in $J^{b}$ as a function of length are included in Fig. S.2, in the Supplementary Information for completeness. Changes in current-direction as indicated by $0-\pi$ oscillations are well known in Josephson junctions with a ferromagnetic weak link Ryazanov et al. (2001); Kontos et al. (2002); Birge and Satchell (2024), but in SNS junctions they are so-far only found when there exists a non-equilibrium electron distribution in N Baselmans et al. (1999). This result therefore indicates that, with regards to generating a Josephson current, strain, or effective SOC, acts on the normal metal in a similar way to a Zeeman interaction in a ferromagnet. A subgap spin Josephson effect has also been suggested in the topologically nontrivial phase of topological Josephson junctions with a magnetic field and SOC in the superconductors Mao and Sun (2022), and a superconducting diode effect has been noted due to a combination of SOC and magnetic field or intrinsically p-wave superconductors Mao et al. (2024). In our case, we show that this current can appear when using conventional s-wave superconductors with bending-strain and a simple normal metal weak link.
In Fig. 5, we show the spin-current $J^{s}$, this time as a function of both strain and position in $N$, alongside the induced spin-magnetization $S^{b}$. A combination of SOC and charge current in normal metals generically leads to spin accumulation Manchon et al. (2015); the charge current shifts the Fermi contour along the $k_{s}$-axis in momentum space, which, in the presence of SOC, creates a finite spin magnetization in the binormal direction. The appearance of a spin-magnetization can therefore be attributed to the Edelstein effect Edelstein (1990). In our numerical scheme, we split the phase difference equally between the two superconductors by fixing the phases in the superconductors. Therefore, the spin accumulation is strongest within the normal metal, as seen in Fig. 5. The diagonalization gives proximity-induced pairings in the normal metal, with a phase gradient $\partial_{s}\phi$. This gradient diminishes away from the interfaces within the superconductors, so the spin magnetization [$S^{b}$] and charge current will eventually decay there.
We see from Fig. 5 that the switch in magnetization can be determined from the $0-\pi$ switching profile of $J^{s}$. The principal reason why no magnetic field or intrinsic p-wave order is required in this case is due to the combination of effective-SOC and interfacial effects on rotating the proximity-induced correlations. It is inherently linked to the thin-film nature of the system, which retains an additional term in the cross-product of the spin vector and the nearest-neighbour vector [Section IV.3] compared with the one-dimensional analysis of most triplet transport analysis to date. When considering two dimensions, we get a $\sigma_{T}\hat{k}_{s}$-term in the $sb$-plane in the presence of a phase gradient. This creates spin-polarized triplets that can engender a magnetization in the binormal direction, and plays the role typically provided by a magnetic field. We therefore see that strain can be an effective generator and controller/switcher of spin-current and magnetization in SNS junctions.
III Discussion
We showed in Section II.2 that bending-strain in an isolated superconducting film induces p-wave triplet pairings throughout the superconductor for uniform curvature. Their magnitude is always around one order lower than the s-wave singlet, and die out as the parent correlation diminishes. Since the superconducting order parameter started out fully s-wave singlet, our self-consistency calculations included only the s-wave gap. We expect that if the p-wave correlations are included in the self-consistency requirement, p-wave pairings may stabilize to form the majority pairing population for increasingly high strain. For non-constant curvatures in the film, the gap would reduce locally in areas with high strain, and the p-wave contribution would increase only locally. We would also expect the sign of the curvature in each arm of the SNS junction to affect the ground state Skarpeid et al. (2024). For an $S$-shaped junction with opposite but equal curvatures in each arm, we would expect the spin magnetization to become vanishingly small as the spin polarization contributions from each interface would cancel Skarpeid et al. (2024).
Bending-strain translates into an effective SOC with symmetry breaking axis in the radial direction. Since the arc length defining the radius of curvature changes for every layer in the film, the total bending-strain of any one sample depends on the film thickness according to Eq. 21. We showed results for thicknesses $N_{n}=(3,6,9)$, and for equal strains these thicknesses have both a different curvature and a different corresponding effective SOC. We provide an overview of their comparative values in table S.1., in the Supplementary Information, were we for convenience also compare the SOC energy with respect to the superconducting gap. When combining the radial bending-strain with interfacial-Rashba SOC in thin-film heterostructures, these two different symmetry-breaking axes combine to give a nuanced mixture of odd- and even-frequency correlations. They can be harnessed in conjunction with the spontaneous charge current that exists in an SNS Josephson junction, giving pure spin-current even at zero phase difference in the junction. The spin current can drive a switchable magnetization in the normal metal. Strain alone may therefore also induce a superconducting diode effect Mao et al. (2024).
In further development of this setup, it would be possible to include intrinsic Dresselhaus SOC in the normal metal, as well as out-of-plane symmetry breaking Rashba SOC induced via a substrate. Furthermore, since the normal metal undergoes no in-situ deformation, it would also be possible to include extrinsic Rashba SOC from gating, giving an additional control over the interconversion mechanisms for the superconducting correlations. While we also restricted the analysis in this study to conventional superconductors, a natural extension would be to consider unconventional materials. For chiral p-wave superconductors for example, bending-strain may split the Fermi surface into two helicity bands Huang et al. (2024), potentially leading to the emergence of chiral Majorana edge states Uchihashi (2021); Gentile et al. (2015); Ahmed et al. (2024). For unconventional or multiband superconductors, modifications to the electron-phonon interaction may play an even bigger role, and strain-deformation of the bands in two-band superconductors could have profound implications for stabilizing superconductivity at high magnetic fields Salamone et al. (2023).
IV Methods
This section summarizes the theoretical and numerical framework used to derive the results presented above. We introduce geometric curvature in superconductors using the Frenet-Serret frame in Section IV.1, and define the continuum Hamiltonian with curvature-induced bending-strain in Section IV.2. We discretize the Hamiltonian to give the tight-binding model in Section IV.3, and in Section IV.4 we provide details of the numerical algorithm and material parameters employed to produce the examples that are analyzed in the main text.
IV.1 Frenet-Serret frame
@@
Here we have introduced the geometric curvature function $\kappa(s)$, which for a uniform curve is simply the inverse of the radius of the curve (see Fig. 1). Using Eq. 13, we can write the basis vectors as derivatives of $\bm{R}(s,n,b)$. We express them as $\bm{e}_{\mu}=\partial_{\mu}\bm{R}=h_{\mu}\hat{e}_{\mu}$, where $h_{\mu}$ are scale factors. The metric tensor is defined as ${\mathcal{G}_{\lambda\mu}=\bm{e}_{\lambda}\cdot\bm{e}_{\mu}}$, and can be written as
@@
$\bm{r}(s)={-}\frac{1}{\kappa}\cos(\kappa s)\>\hat{\bm{e}}_{x}+\frac{1}{\kappa}\sin(\kappa s)\>\hat{\bm{e}}_{y}.$ (15)
The orthogonal curvilinear unit vectors become $\hat{\mathcal{T}}(s)=\partial_{s}\bm{r}(s)$, $\hat{\mathcal{N}}(s)={-}\partial_{s}\hat{\mathcal{T}}(s)/|\partial_{s}\hat{\mathcal{T}}(s)|$ and ^$\hat{\mathcal{B}}(s)=\hat{\mathcal{T}}(s)\times\hat{\mathcal{N}}(s)$. For the circular parametrization [Eq. 15] this gives
$\begin{split}\hat{\mathcal{T}}(s)&=\sin(\kappa s)\>\hat{\bm{e}}_{x}+\cos(\kappa s)\>\hat{\bm{e}}_{y},\\ \hat{\mathcal{N}}(s)&=-\cos(\kappa s)\>\hat{\bm{e}}_{x}+\sin(\kappa s)\>\hat{\bm{e}}_{y},\\ \hat{\mathcal{B}}(s)&=\hat{\bm{e}}_{z}.\end{split}$ (16)
IV.2 Continuum Hamiltonian and strain-induced spin-orbit coupling